OpenStudy (anonymous):

2/3 of x equals 3/4 of x , x= ?

OpenStudy (anonymous):

2x/3=3x/4, then ??

OpenStudy (anonymous):

8x=9x , then ???

OpenStudy (anonymous):

How did you get 8x=9x?

OpenStudy (anonymous):

3* 3x= 4*2x

OpenStudy (anonymous):

Ok good now for the next part we can just use subtraction

OpenStudy (anonymous):

???

OpenStudy (anonymous):

Are you still lost?

OpenStudy (anonymous):

:( yes how to subtract here???

OpenStudy (anonymous):

By the way do you know the solution so that we can check?

OpenStudy (anonymous):

noo, here it comares 4/5 of x with 0

OpenStudy (anonymous):

compares

OpenStudy (anonymous):

What do you mean compares

OpenStudy (anonymous):

which is greater 4/5 of x or 0, if i cold find x acc. to 2/3 of x equals 3/4 of x , x= ?, so I could copare them

OpenStudy (anonymous):

Alright well going back to your quesiton when we get the 8x=9x we can just just subtract 8x from both sides